OpenStudy (anonymous):

Really need the answer! Find parametric equations for the curve y = x + 1, x ≥ 0, by using as parameter the distance of the point (x,y) from the origin.

OpenStudy (anonymous):

i m seeing a page and there they says that we cant because we need to express x in terms of the parameter but in this case d=sqrt(x^2 +y^2) => d= sqrt(x^2 +(x+1)^2) if we want to solve for x we will get 2 values of (d) so we cant http://www.phengkimving.com/calc_of_one_real_var/13_plane_curves/13_01_param_curves/13_01_01_param_curves.htm

OpenStudy (anonymous):

i dont have a lot that is why i m searching for that subject could u solve that problem?

OpenStudy (anonymous):

but the parameter must be the distance to the origin that is d=sqrt(x^2 +y^2)

OpenStudy (anonymous):

what u did there is just get parametric equations with the parameter (t) but in this case the parameter must be (d)

OpenStudy (primeralph):

got to go. ...